# On the Geometry of Semantics in Next-token Prediction

### v1

@@ 5.1 Setup
Visuals. To interpret the semantic information, we visualize the top $40$ words/contexts associated with each concept. “Top” here is measured with respect to the typicality score in Defn. 1. For visualization, we use word clouds where size is proportional to its typicality score. We often use the pound sign (#) to indicate token start.
5.2 Findings
Individual concepts are not interpretable. We have seen in Sec. 4 that not all concepts individually correspond to linguistically interpretable factors. For completeness, we verify this conclusion for the Simplified TinyStories/WikiText datasets in Fig. 14 in the appendix. This observation is well-reported in the literature, e.g., Chersoni et al. [2021], and is also related to the recently popularized superposition hypothesis Elhage et al. [2022].
Combining concepts with signature-configurations reveals semantics. We demonstrate that combining non-interpretable concepts can reveal human-interpretable semantic information. Here, we show this by implementing the idea of orthant-based clustering of Sec. 4.3. Specifically, we combine $p=5$ to $7$ concepts in successive order starting from the most important one, i.e. $\mathcal{K}=\{1,\ldots,p\}$. Fig. 4 shows human-interpretable semantics identified by orthant-based clustering in Simplified TinyStories and WikiText-2 datasets. In Fig. 11, we show how orthant-based clustering progressively reveals hierarchical semantic structure as we combine more concept dimensions, illustrated through the example of verb forms organized by grammatical distinctions. Starting with a single dimension ($\mathcal{K}=\{1\}$), the categories remain broad and difficult to interpret. However, as we incorporate additional dimensions, increasingly refined semantic categories emerge, ultimately distinguishing between different verb forms such as past/present tense, and present continuous forms.
@@ 5.2 Findings
Figure 4: Semantic clusters based on signature configurations of Sec. 4.3. Top: Simplified TinyStories dataset showing Past-tense verbs, Present-tense verbs, Prepositions, and Proper names. Bottom: Simplified Wikitext dataset showing Months, Numbers, Units, and Prepositions/Conjunctions. Each configuration is shown in its title.
Past-tense verbs Refer to caption
Proper names Refer to caption
Numbers Refer to caption
Conjunctions Refer to caption
$k$-means spectral clustering. As a complementary approach to our orthant-based clustering, we also investigate k-means spectral clustering. Specifically, we perform k-means clustering according to pairwise distances of analyzer vectors with their top $p=\log_{2}(k)$ dimensions (corresponding to top-$p$ most important concepts). Vanilla $k$-means spectral clustering selects $k$ principal dimensions. The modification here is inspired by the orthant-based clustering for which $p$ concepts correspond to $2^{p}$ orthants. Fig. 5 shows semantic meaning in a subset of the recovered clusters with $k=32$ and $k=64$ for Simplified TinyStories and WikiText-2, respectively; see Sec. F.5 for full results. See also discussion in Sec. A.3.
Relation of words and context semantics. Fig. 2A demonstrates a relationship between context embeddings and word embeddings that share semantic features: contexts and words representing the same semantic concept (e.g., plant or animal) cluster proximally in the embedding space. This pattern extends to the richer datasets we study here. To verify this, we perform orthant-based clustering of contexts and words based on the same signature configuration of a subset of concept dimensions. Fig. 6 shows this for Simplified TinyStories: contexts containing modal constructions like "want/like/love/have to" correspond strongly with word clusters containing infinitive verbs. Similarly, contexts with intensifier patterns "was so/very" align with clusters of descriptive adjectives. This relation reflects a natural dual semantic interpretation of contexts: they can be understood either through the direct semantic meaning of their constituent words or through the semantic properties of words that typically follow them. The latter interpretation aligns naturally with the predictive nature of NTP, where contexts acquire meaning partly from what they predict.
  Refer to caption Refer to caption
Figure 6: Context-word cluster pairs sharing identical sign patterns across top 5 concepts. Left and right pair illustrates the “verb+to infinitive” and “was so/very + adj” structures, confirming semantic alignment between contexts and their in-support words.
Semantics in pretrained models. We find interpretable semantics from simplified datasets are primarily grammatical or syntactic due to limited sample size and context simplicity. While these features make analytical computations feasible (e.g., computing/factorizing $\widetilde{\bm{S}}$), they compromise the richness of sparsity patterns and, in turn, the emergence of semantics. To verify that richer sparsity patterns yield richer semantics, we additionally experiment with pre-trained GPT-2. See Fig. 7 and Sec. F.2 in appendix for additional results.
@@
Figure 7: Semantics identified by orthant-based clustering on GPT2’s word embeddings.
6 Conclusion and Limitations
While continuing to advance LLMs’ capabilities, understanding their inner workings is crucial. We study how NTP optimization guides the geometry of word and context representations to organize around latent semantic information. Without explicitly computing it, the model encodes such information in the SVD factors of an implicit data-sparsity matrix. As an early contribution in this direction, our analysis motivates several future directions: (1) While the assumption $d\geq V$ allows for rich geometric structures, a formal investigation is needed for the $d<V$ case prevalent in practice (see Sec. A.1). (2) Our analysis assumes sufficient expressivity and training time, which may become increasingly relevant with improved compute or in data-limited regimes, but relaxing them or experimentally validating their applicability in the spirit of our GPT-2 experiments (e.g., in the spirit of Wu and Papyan [2024]) is important. (3) Understanding how transformer architectures specifically fit into our analytical framework remains an open challenge. (4) Some conclusions rely on square loss and our work highlights a significant gap in understanding CE loss dynamics, even in simple bilinear models. (5) While our approach differs from concurrent efforts to explain semantic emergence in LLMs via probing or sparse-dictionary learning, future work could explore connections and potential applications to model improvement techniques.
@@ 7 Acknowledgments
Appendix A Additional Discussions
A.1 On the Assumption $d\geq V$
Our analysis has assumed the embedding dimension $d$ is at least as large as the rank of the data-sparsity matrix (guaranteed when $d\geq V$), allowing word and context embeddings to form geometric structures representing all concepts encoded in the matrix’s singular factors Zhao et al. [2024]. As noted in Zhao et al. [2024], while this assumption differs from current practice in LLMs, the geometry of embeddings remains rich in this setting. Importantly, this assumption is less restrictive than requiring $d>C$ in one-hot classification with $C$ classes. There, due to collapse of embeddings from the same class, $d>C$ effectively requires the dimension to exceed the number of training examples. In contrast, for NTP, the number of contexts $m$ can be (and typically is) much larger than $d$, allowing for rich geometric arrangements of context embeddings in the lower-dimensional space and non-trivial emergence of concepts, which is confirmed throughout our case studies.
That said, the current trend of modern language models typically employing $d<V$, raises the question: Which concepts are prioritized when the model cannot capture all of them? We hypothesize that during NTP training, models learn to represent the $d$ most significant concepts, corresponding to the largest singular values of the data-sparsity matrix. While this requires a separate study that is beyond our current scope, our preliminary investigation using NTP-UFM with square-loss on a synthetic dataset supports this hypothesis. Fig. 19 shows the learned logit matrix’s singular values converging to the $d$ largest singular values of the data-sparsity matrix during training. Following Sec. 4.4, we used square loss rather than CE to ensure bounded singular values. While a rigorous analysis for $d<V$ remains future work, since meaningful semantic categories emerge from combinations of concepts, even a reduced set of core concepts could enable rich semantic representations through the orthant clustering mechanism of Sec. 4.3.
A.2 Role of Autoregression
We have shown that concepts emerge during NTP training as principal directions of $\widetilde{\bm{S}}$, a centered version of the support matrix $\bm{S}$. Each column $j$ of $\bm{S}$ corresponds to a context and can be viewed as a binary (multi-)label vector $\bm{s}_{j}$, where entries of $1$ indicate tokens that appear as next-tokens for that context in the training data. While concepts are determined by the principal directions of this label matrix, this might seem limiting: contexts can relate not only through shared next-tokens but also through their intrinsic structure (e.g., overlapping constituent tokens).
The autoregressive nature of training naturally captures these structural relationships as follows. As the model processes labels for progressively longer contexts, each context $(z_{1},\ldots,z_{t})$ contributes to concept formation both through its distribution over next-tokens $z_{t+1}$ and through the labels of its shorter subsequences $(z_{1},\ldots,z_{t-1})$, $(z_{1},\ldots,z_{t-2})$, etc.—each representing distinct columns in $\bm{S}$. This overlapping window structure produces fine-grained label information, creating rich sparsity patterns in $\bm{S}$ (and consequently in $\widetilde{\bm{S}}$) that yield nontrivial concepts with varying significance, as reflected in word and context embeddings.
A.3 Relationship to Spectral Clustering
The orthant-based clustering approach introduced in this work shares foundational similarities with spectral clustering while providing distinct advantages for linguistic interpretation. Both methods operate in reduced-dimensional spaces and leverage embedding geometry to identify meaningful clusters. First, while both approaches leverage singular factors, the orthant-based method provides a more direct link (in view of concept definition in Sec. 4) between NTP and the emergence of semantic structure by explicitly interpreting the sign patterns of SVD dimensions as semantic indicators.
An interesting property of our method is its selective nature: while $p$ dimensions generate $2^{p}$ potential orthants, only a subset contains semantically coherent word groupings. This selectivity reflects linguistic reality, where not all mathematical combinations yield meaningful semantic categories. In contrast, spectral clustering typically partitions the space completely, which may not always align with natural semantic boundaries in language. Finally, the orthant-based method offers semantic interpretation of membership and typicality, and dimensionality control through parameter $p$. This aligns with language model training dynamics, as the most important singular value components are learned first during training, making the resulting clusters reflective of the model’s knowledge acquisition process. To this respect, note that unlike orthant-based clustering producing a potential maximum of $2^{p}$ clusters when selecting $p$ top concepts, typical clustering algorithms such as vanilla $k$-means on principal components would result in $p$-clusters with the same number of selected concepts. Thus, when we evaluate spectral methods in our experiments, we choose $k$ on the order of $2^{p}$ for the value of $p$ that we find orthant-based clustering reveals semantically informative orthants.
A.4 Neural Collapse Meets Semantics
The emergence of concepts with rich semantic meanings stems from the interplay of labels (encoded in $\bm{S}$ or its centered version $\widetilde{\bm{S}}$) across different contexts. Consider the minimal label richness case: each context has exactly one next-token, with contexts distributed equally across the vocabulary. Here, $\bm{S}$ can be rearranged as $\mathds{I}_{V}\otimes\mathds{1}_{m/V}^{\top}$, yielding trivial concepts where all singular directions contribute equally. This setting parallels standard balanced one-hot classification studied in neural collapse literature, where last-layer embeddings and weights form highly symmetric aligned structures Papyan et al. [2020a], reflecting the symmetric nature of $\bm{S}$ where labels induce no interesting conceptual structure.
However, such balanced settings never occur in autoregressive NTP, which as discussed in Sec. A.2, produces rich label formations. In these richer settings, $\bm{S}$ induces a geometry of concepts that provides semantic interpretation to the embedding geometry studied in the neural collapse literature. To illustrate this new perspective, we demonstrate that meaningful concepts emerge even in one-hot classification—the traditional focus of neural collapse literature—given minimal deviation from perfect balance. Specifically, consider STEP imbalances with ratio $R$: $V/2$ majority classes each have $R>1$ times more samples than minority classes.
Here, Thrampoulidis et al. [2022] shows the learned logit matrix converges to $\widetilde{\bm{S}}$, whose singular values exhibit a three-tier structure:
$\sigma_{1}=\ldots=\sigma_{V/2-1}=\sqrt{R}>\sigma_{V/2}=\sqrt{(R+1)/2}>\sigma_{V/2+1}=\ldots=\sigma_{V}=1.$ The left singular vectors matrix $\bm{U}$ takes a sparse block form:
where $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ is an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$. [^7] For concreteness, $\mathbb{F}$ can be constructed using the discrete cosine transform matrix, excluding the constant column: $\mathbb{F}[i,j]=\sqrt{\frac{4}{V}}\cdot\cos\left(\frac{\pi(2i-1)j}{V}\right)$ for $i\in[V/2]$, $j\in[V/2-1]$ The structure of $\bm{U}$ reveals three distinct types of concepts, corresponding to the three tiers of singular values: (1) First $V/2-1$ columns (non-zero only for majority classes) represent distinctions among majority classes; (2) Middle column (singular value $\sqrt{(R+1)/2}$) has opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy; (3) Last $V/2-1$ columns (non-zero only for minority classes) capture distinctions among minority classes. This hierarchical structure shows the network learns concepts in order: first majority class distinctions, then the majority-minority split, and finally minority class differences. See Fig. 13 for a visualization of $\bm{U}$ and Fig. 9 for an experiment that confirms the above semantic interpretation of concepts (columns of $\bm{U}$).
Figure 9: Experimental illustration of the fact that important concepts are learned first. Confusion matrix evolution during training of a 3-layer convolutional network on an imbalanced MNIST data. The matrices show five snapshots at different training steps (0, 13, 30, 55, 65) for four classes: two majority classes (Maj1, Maj2) with 100 samples each and two minority classes (Min1, Min2) with 10 samples each. Training progresses from left to right: initially classifying all data as Maj1 (Step 0), then correctly identifying majority classes while misclassifying minority classes (Step 13), gradually improving minority class recognition (Step 30), confusion between minority classes only (Step 55), and finally achieving perfect classification by Step 60.
Appendix B From Sparsity Language Pattern to Concepts Geometry

### v2

@@ 4.1 Methodology
To interpret semantics, we visualize top 40 words/contexts associated with each orthant. “Association” and “top” are evaluated per Defn. 1. We use word clouds with size proportional to typicality score. Symbol # indicates token start.
4.2 Results
Individual concepts are not interpretable. We verify that not all individual concepts correspond to linguistically interpretable categories (Sec. 3.1). Fig. 12 in App. demonstrates this for Simplified TinyStories and WikiText. This aligns with findings in the recent literature [Chersoni et al., 2021, Piantadosi et al., 2024, Elhage et al., 2022].
@@ 4.2 Results
Figure 3: Semantics from orthant-based clustering on $\widetilde{\bm{S}}$ in Simplified TinyStories (Top) and Simplified Wikitext (Bottom). Semantic interpretations and combination configurations are shown in the title.
@@
Figure 4: Semantics identified by orthant-based clustering on GPT-2’s word embeddings.
Combining concepts with sign configurations reveals semantics. We verify that combining (non-interpretable) individual concepts reveals interpretable semantics via orthant-based clustering (Sec. 3.2). Fig. 3 shows human-interpretable semantic categories identified by orthant-based clustering in Simplified TinyStories and WikiText. In Fig. 11 in Appendix, we further demonstrate the hierarchical semantic structure that emerges as we combine more concepts, illustrated through verb forms organized by grammatical properties. Starting with only the top concept ($k=1$), the categories remain broad and difficult to interpret. However, as we incorporate additional concepts, increasingly refined semantic categories emerge, ultimately distinguishing between different verb forms (past tense, present tense, and present continuous).
Orthant-based clustering can be viewed as instance of spectral clustering Von Luxburg [2007]. We thus also explore k-means spectral clustering over concept subspaces in App. D.8. While this also reveals interpretable categories, orthant-based clustering offers advantages: (1) It provides a direct link between optimization and categories by explicitly interpreting sign patterns of SVD dimensions as semantic indicators. (2) It has built-in selectivity: not all $2^{p}$ potential orthants yield semantically coherent groupings—only a meaningful subset does. (3) This control over $p$ enables semantic interpretation of membership and typicality, and allows exploration of different semantic granularities.
Figure 5: Semantics from transformers of different depth, both trained on Simplified TinyStories.
Figure 6: Context/word semantics share identical configurations. Left and right pairs illustrate the “verb+to infinitive” and “was so/very + adj” structures, confirming semantic alignment between contexts and their in-support words.
Relation of words and context semantics. Recall from Secs. 3.1,3.2 that semantic clustering of both word and context embeddings is guided by the SVD of the same data-sparsity matrix $\widetilde{\bm{S}}$. This coupling predicts that words and contexts sharing semantic properties should cluster together in embedding space. Fig. 2A illustrated this phenomenon in a simple setting, where plant-related words and contexts clustered separately from animal-related ones. We now demonstrate this in more complex datasets: We apply orthant-based clustering to words and contexts using the same signature pattern across a subset of concept dimensions extracted from Simplified TinyStories. Fig. 6 shows that contexts with modals like “want/like/love/have to” align with clusters of infinitive verbs; intensifiers like “was so/very” correspond to descriptive adjectives. This duality reflects how contexts encode meaning both via their surface words and via the semantics of likely next tokens—a view naturally supported by the predictive nature of NTP.
Semantics in pretrained models. We find that simplified datasets primarily reveal grammatical/syntactic structure due to limited size and context diversity. To verify that richer sparsity induces richer semantics, we here analyze pretrained models. In GPT-2 (Fig. 4), orthant-based clustering recovers categories such as medical terms, emotional tone, entertainment, and political figures. In BERT (Fig. 14 in App.), we find clusters reflecting numerical patterns (e.g., 3-digit numbers), action verbs, administrative phrases, and morphological forms. Applying the same method to Qwen-7B yields UI phrases, programming keywords, numeric tokens, functional Chinese expressions, and so on. Qwen requires combining more concepts (7–11 vs. 5–7 for GPT-2) to isolate coherent groups (Table 1 in App.), suggesting more distributed semantic encoding, potentially reflecting its multilingual training.
Consistent structure across models. Our analysis predicts that, as long as an architecture is sufficiently expressive for the UFM assumption to hold, the geometry of learned semantics should depend on data patterns rather than specific architectural details. We verify this by training two transformers with different depths (6 vs. 12 layers) on the same Simplified TinyStories dataset. After training, we extract the decoder matrix $\bm{W}$} from each model and compute its top three left singular vectors. Applying orthant-based clustering, both models yield nearly identical semantic groupings (Fig. 5).
5 Related Work
We bring together, in the context of NTP-trained language models, two classical ideas: (1) semantics emerge from linear relationships between vector representations of textual inputs, and (2) learned-representations geometries are tied to language statistics via count matrices. App. B for details.
Extracting semantics from distributed representations was popularized through word embedding research Mikolov et al. [2013a], with approaches including semantic axes An et al. [2018], Fast et al. [2016] and word analogies Mikolov et al. [2013b], Levy et al. [2015]. In LLMs, work has focused on dictionary learning Turner et al. [2023], Li et al. [2024], geometric clustering Coenen et al. [2019], Hewitt and Liang [2019], and linear probing Hewitt and Manning [2019], Marks and Tegmark [2023]. Unlike post-hoc analyses, we analyze, from an optimization perspective, how semantics emerge as a consequence of training. While Park et al. [2023, 2024] have explored similar questions, here we directly analyze NTP optimization, rather than relying on a generative model for concept generation and its link to words/contexts.
Classical work has shown that applying factorization methods to co-occurrence matrices conveys semantic information Landauer [1997], Levy et al. [2015], with various hand-engineered choices for both the matrix representation (e.g., word-document, word-word, PMI, PPMI) and factorization technique (SVD, NNMF Lee and Seung [1999], Ding et al. [2006], sparse dictionary learning Murphy et al. [2012], Faruqui et al. [2015]). We demonstrate, for the first-time, that gradient-based optimization itself determines both the matrix representation ($\widetilde{\bm{S}}$) and factorization method (SVD) used to encode semantics.
Our work connects NTP-UFM to the closed-form framework of Saxe et al. [2013, 2019] for linear neural networks. While their work focuses on general cognitive development with orthogonal inputs, we provide a new practical instantiation for language modeling optimization and validate theory findings on real data and architectures.
6 Conclusion
As LLM capabilities continue to advance, understanding their inner workings is crucial. We study how language-modeling optimization guides the geometry of word and context representations to organize around latent semantic information. Without explicitly computing it, the model encodes such information in the SVD factors of an implicit data-sparsity matrix. As an early contribution in this direction, our analysis motivates several future directions, some of which we outline in Sec. 7.
7 Limitations
(1) While the assumption $d\geq V$ allows already for rich geometric structures, most practical models use $d<V$. We hypothesize that during NTP training, models learn to represent the $d$ most significant concepts, corresponding to the largest singular values of $\widetilde{\bm{S}}$. While NTP-UFM offers preliminary support of this hypothesis, a rigorous theoretical and experimental analysis requires separate future investigation. Encouragingly, since semantic categories arise from concept combinations (Sec. 3.2), even a reduced set of core concepts could enable rich semantic representations. (2) Our analysis assumes sufficient expressivity and training time, which may become increasingly relevant with improved compute or in data-limited regimes, but relaxing them or experimentally validating their applicability in the spirit of our GPT-2/Qwen experiments is important. (3) Existing theoretical work on transformer optimization dynamics is limited to shallow models (e.g. one-layer) and overly strict data-generation assumptions that limit capturing meaningful semantic information. On the other hand, the abstraction of unconstrained features offers mathematical tractability, already useful insights, and has the attribute of leading to predictions insensitive to the specifics of the architecture (Sec. 4.2). Future work can explore merging the two paradigms of analysis. (4) While our approach differs from concurrent efforts to explain semantic emergence in LLMs via probing or sparse-dictionary learning, future work could explore connections and potential applications to model improvement techniques.
@@ 8 Acknowledgments
Appendix A Paper Organization
The main paper is organized as follows: Sec. 2 introduces background on our analytical model and the geometry of embeddings. Our main analytical contributions and minimal synthetic experiments are presented in Sec. 3, with subsections 3.13.3 each detailing one of the three contributions (C1-C3). Sec. 4 provides qualitative studies on real-world data and pretrained models that validate our findings. Sec. 5 contextualizes our work within the broader literature, with comparisons to closely related work integrated throughout the paper. We conclude in Sec. 6 and discuss limitations and future directions in Sec. 7.
The appendix is organized as follows: Sec. B provides an extended discussion of related work (a condensed version appears in the main paper). Sec. C includes details on the formalization of concept representations discussed in Sec. 3.1 and the proof of Theorem 1 from Sec. 3.3. Additional experimental results deferred due to space constraints appear in Sec. D.
Appendix B Extended Discussion on Related Work
This section provides an extended discussion of the related work summarized in Sec. 5.
Semantics from Distributed Representations. The idea of extracting semantics from distributed representations, while rooted in earlier works Bengio et al. [2000], was popularized through word embedding research Mikolov et al. [2013a]. Two classically popular approaches include: (i) constructing semantic axes from centroids of pole words An et al. [2018], Fast et al. [2016], evolving from lexicon-based sentiment analysis Taboada et al. [2011], Hu and Liu [2004], and (ii) investigating semantic patterns through “word analogies,” which reveal that semantic relationships manifest as linear directions in embedding space Mikolov et al. [2013b], Levy et al. [2015], Drozd et al. [2016], Ethayarajh et al. [2019], Allen and Hospedales [2019]. In the context of modern LLMs, word analogies have been empirically observed Rezaee and Camacho-Collados [2022], Wijesiriwardene et al. [2024]. However, recent focus has shifted to dictionary learning techniques for distilling interpretable semantics, e.g., Turner et al. [2023], Li et al. [2024]. Alternative approaches include geometric studies showing how linguistic features cluster in embedding spaces Coenen et al. [2019], Hewitt and Liang [2019]. Simultaneously, supervised methods such as linear probing have been employed to extract semantic and syntactic information from modern LLMs Alain and Bengio [2017], Hewitt and Manning [2019], Marks and Tegmark [2023], relying on the same principle of linear representation found in word analogies. Unlike these works that empirically analyze trained models post-hoc, our goal is not to introduce a state-of-the-art method for extracting semantics, but rather to understand, from an optimization viewpoint, how semantics emerge as a consequence of next-token prediction. While Park et al. [2023, 2024] have explored similar questions, here we directly analyze NTP optimization, rather than relying on a generative model for concept generation and its link to words/contexts.
Geometry from Co-occurrence Statistics and Matrix Factorization. The idea that factorizations of co-occurrence matrices convey latent semantic information dates back to latent semantic analysis of word-document co-occurrence matrices Landauer [1997], Deerwester et al. [1990] and later to pointwise mutual information matrices of word-word co-occurrences Levy and Goldberg [2014], Levy et al. [2015]. Beyond SVD, alternative factorization techniques such as non-negative matrix factorization Lee and Seung [1999], Ding et al. [2006] and sparse dictionary learning Murphy et al. [2012], Faruqui et al. [2015] have been applied to various transformations of co-occurrence data, including those addressing PMI sparsity (e.g., PPMI) or using probability smoothing Levy et al. [2015].
More broadly, vector space representations emerge from two approaches: (i) traditional count-based models using co-occurrence matrices, and (ii) prediction-based models that optimize token prediction objectives like Skip-gram or NTP. Levy and Goldberg [2014], Pennington et al. [2014] demonstrate that these approaches are fundamentally similar, as both probe underlying corpus co-occurrence statistics. Zhao et al. [2024] have extended this insight to causal NTP models with a key distinction: the co-occurrence matrix between contexts and words is inherently sparse, unlike the dense word-word co-occurrences in prior work. This distinction makes relevant recent advances in implicit regularization Soudry et al. [2018] and neural collapse geometries Papyan et al. [2020b], Mixon et al. [2020].
Relation to [Zhao et al., 2024]. Our work extends Zhao et al. [2024] by addressing the three open questions Q1-Q3 discussed in our Introduction. In doing so, we provide crucial semantic interpretation to prior NC-based representation analyses (including those for imbalanced one-hot settings Fang et al. [2021], Thrampoulidis et al. [2022a]). Furthermore, by connecting our model to the closed-form analysis of linear neural network learning dynamics Saxe et al. [2013], we rigorously demonstrate how singular values govern the order of concept acquisition during training.mine the learning order during training.
Relation to Saxe et al. [2019]. Finally, while drawing conceptual similarities to [Saxe et al., 2019], our work differs in its motivation, interpretation, and scope of results. Their focus is on general cognitive development using an abstract two-layer linear neural network with orthogonal inputs. In contrast, we specifically analyze how semantic and grammatical concepts emerge from natural language data through NTP optimization. Methodologically, we realize a connection and apply their framework to the NTP-UFM (which we interpret as a linear two-layer network where the input dimension equals the number of input examples), a tractable model for expressive large architectures. This provides a practical instantiation for their otherwise restrictive orthogonal-input assumptions. We also corroborate our findings with experiments on real datasets and architectures, a crucial distinction from their purely theoretical work.
Appendix C Derivations and Proofs
C.1 Concepts in Embedding Space